ADNAN MUHAMMED

Grade 12,

Prove that the given straight lines always meet on a hyperbola.Given lines are : x\a - y\b = m and x\a + y\b = 1\m.

Prove that the given straight lines always meet on a hyperbola.Given lines are : x\a - y\b = m and x\a + y\b = 1\m.

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
When we write a combined equation of these straight lines,
(x/a - y/b) (x/a + y/b) = m* 1/m
x²/a²  - y²/b² = 1
This is the equation of hyperbola.
Hence proved
 
Regards
Arun (askIITians forum expert)